OpenStudy (anonymous):

Determine the solution to f(x) = g(x) using the following system of equations: f(x) = 3x + 12 g(x) = −9.5x − 13

OpenStudy (campbell_st):

as it says... let f(x) = g(x) so 3x + 12 = 9.5x -13 and solve for x

OpenStudy (anonymous):

x=-2 x=-3 x=4 x=5

OpenStudy (anonymous):

those are the choices

OpenStudy (anonymous):

so do i swap f(x) and g(x) or do i get rid of f(x) and replace it with g(x)?

OpenStudy (campbell_st):

well start by adding 13 to both sides of the equation 3x + 12 = 9.5x - 13 + 13 + 13 -------------------- 3x + 25 = 9.5x -3x -3x subtract 3x from both sides of the equation --------------- 25 = 6.5x now solve for x..... you may like to check you have f(x) and g(x) correct as the solution is a decimal...

OpenStudy (anonymous):

i got 3.84615384615

OpenStudy (anonymous):

so do i just round up to 4?

OpenStudy (anonymous):

or keep it at 3

OpenStudy (campbell_st):

that's why I asked if you had entered f(x) and g(x) correctly

OpenStudy (anonymous):

i copy and paste it ill check still tho

OpenStudy (campbell_st):

if you round the answer, it goes to 4

OpenStudy (anonymous):

ik that and im guessing thats what i need to do because i checked and its right

OpenStudy (campbell_st):

ok... good luck

OpenStudy (anonymous):

wait you for got to put 9.5x as a negative when you solved it

OpenStudy (anonymous):

i just noticed

OpenStudy (anonymous):

ohh wait its still 6.5 because you have to combine them :/

OpenStudy (anonymous):

-4 isnt a choice...

OpenStudy (anonymous):

its either b or c

OpenStudy (campbell_st):

hold on I'll redo it correctly and now it makes more sense 3x + 12 = -9.5x - 13 +13 +13 add 13 ------------------ 3x + 25 = -9.5x -3x -3x subtract 3x from both sides --------------- 25 = -12.5x now the solution should make sense

OpenStudy (anonymous):

so its a?

OpenStudy (anonymous):

YAY A WAS RIGHT I GOT A 100 ON HOMEWORK!